The length of a rectangular is 10 feet more than twice its width. the perimeter of the rectangule is 170feet. find the dimension of the rectangle
please reply me the answer in detail steps. thanks!

Answer by Paul(988)

Length = 10+2x
width = x
Formula = 2(l+w)=P

2(x+2x+10)=170
6x+20=170
6x=150
x = 25

25(2)+10= 60

Hence the length is 60ft and the width is 25ft.
Paul.